If the square of a number decreased by the number is 56, then what is the numbers?
Answer by mananth(16946)

let number be x
square of number = x^2
decreased by x
x^2-x
x^2-x=56
x^2-x-56=0
x^2-8x+7x-56=0
x(x-8)+7(x-8)=0
(x-8)(x+70=0
x = 8 OR -7
...
m.ananth@hotmail.ca